# Particles before symmetry

### v4

@@ 5.1 Slack between symmetry and geometry: examples and lessons
The first step to describe the differences between the PFB-POV and the VB-POV is to recall how the two points of view differ in their basic ingredients, as illustrated in Figures 1 and 2, respectively. In the VB-POV, one can only define groups $G$ via products of automorphism groups, $\mathsf{Aut}(V_{n})$, where here $n$ indexes the family of fundamental vector bundles (of course, one is under no obligation to invoke even the latter groups; it is not part of the ground level of explanation). One can reconstruct certain $(P,G,M)$, by using $L_{\text{\tiny{adm}}}(E_{n})$, namely, using the bundle of admissible frames for $E_{n}$ (see Section 2.1). This reconstruction fixes $G\simeq\prod_{n}\mathsf{Aut}(V_{n})$ and $\rho=\prod\rho_{n}^{o}$, where $\rho_{n}^{o}$ are fundamental representations (one can also use symmetrisations and direct products to recover different representations).
By contrast, in the PFB-POV there are in principle three separate ingredients: not only $G$ (or $P$) and $V$, but also the representation $\rho$. $P$ is just a manifold admitting the free and proper action of a Lie group, and an associated vector bundle can be built by using any representation of that group on any vector space $V$. There is no requirement that $G$ transparently reflect the structure of $V$. Thought in this way, the question is not whether one can reconstruct a given $G$ as the automorphism of some $V$, but rather why should the geometry of $V$ reflect the, in principle independent group that acts on $P$?
In the equivalence relation that defines associated bundles, (2.1), $\rho(G)$ need not coincide with the automorphism group of the typical fibre $V$. The only requirement is that the gauge group preserve the structure of the vector spaces:
@@
To sum up the contrast: in the spacetime case we have an identification,
$SO(3,1)=G\simeq\rho(G)\simeq\mathsf{Aut}(T_{x}M).$ (5.2)
And so it is for principal bundles recovered from the VB-POV,
@@
Here I won’t be so interested in the full recovery of the geometry from the group. [^27] In a strict sense, to fully recover that geometry, the action might be asked to be faithful and transitive on each fibre: it should be able to carry any element of a given fibre to any other. For instance, in ordinary differential geometry, consider the so-called surfaces of revolution. These surfaces possess rotational symmetry, yet that symmetry does not determine their full geometry—think of the diversity of shapes (or vases) that share the same axis of rotation. On the broader question of recovering the geometry of a vector space from the action of subgroups of $GL(V)$, or, more broadly, the geometry of a space from the action of pseudo-groups, see \@BBOPcitep\@BAP\@BBN(Wallace, 2019; Barrett & Manchak, 2024)\@BBCP . It suffices for my purpose to say that a group reflects the geometry when both conditions of (5.3) are met.
Let us first recall from Section 2.1 two sets of relations: (i) between $P$ as an arbitrary principal fiber bundle with an arbitrary representation $\rho$ on the typical fiber $V$ of $E$; and $P^{\prime}:=L_{\text{\tiny{adm}}}(E)$—the bundle of admissible frames of $E$, defined solely in terms of the structure of EE—and (ii) between $E$ and the associated vector bundle $E^{\prime}:=P\times_{\rho}V$.
$P^{\prime}=L_{\text{\tiny{adm}}}(E)$ is a principal fiber bundle over $M$ with structure group $G^{\prime}\simeq\mathsf{Aut}(V)$. Thus, when (5.3) holds, $G\simeq\mathsf{Aut}(V)\simeq G^{\prime}$ and $P$ is at least locally isomorphic to $P^{\prime}$, since they have isomorphic fibers and base space, and so there are open subsets $U\subset M$ in which $P_{|U}\simeq U\times G\simeq P^{\prime}_{|U}$ (the isomorphism is a fiber-respecting diffeomorphism). In this case $E\simeq E^{\prime}=P\times_{\rho_{o}}V$ for $\rho_{o}$ the fundamental representation (see \@BBOPcitep\@BAP\@BBN(Weatherall, 2016)\@BBCP for details). [^28] The image of any two representations $\rho,\rho^{\prime}$ for which (5.3) holds are isomorphic to $\mathsf{Aut}(V)$ and thus they must be related by a similarity transformation: $\rho^{\prime}=S^{-1}\rho S$, for $S\in GL(V)$. Therefore the map $(p,v)\mapsto(p,Sv)$ (5.4) gives rise to a local isomorphism between $P\times_{\rho_{o}}V$ and $P\times_{\rho}V$. Thus, when (5.3) holds, we get local isomorphisms $P\simeq L_{\text{\tiny{adm}}}(E)$ and $E\simeq P\times_{\rho}V$—but this second condition always holds, as I show below. And I won’t here concern myself with global isomorphisms; my hunch is that their existence will depend on the topology of $M$.
Conversely, in case $(P,M,G)$ is such that $G\not\simeq\mathsf{Aut}(V)$, $P$ is no longer isomorphic to $P^{\prime}=L_{\text{\tiny{adm}}}(E)$, since $G^{\prime}\simeq\mathsf{Aut}(V)$. So one shouldn’t think of $P$ as a frame bundle. But it is still the case that, $E^{\prime}:=P\times_{\rho}V$ is well-defined and, at least locally, $E\simeq E^{\prime}$. This can be established from (2.3), by which we can describe any section of $P\times_{\rho}V$ over $U\subset M$ using a local $V$-valued field on $M$ by first fixing a local section of $P$.Since a section $\sigma:U\rightarrow P$ is an embedding, its image is diffeomorphic to $U$. Thus locally $P_{|U}\times_{\rho}V\simeq U\times V\simeq E_{|U}$. Since the structure of $E$ is inherited from the structure in $V$, we obtain a local isomorphism $E\simeq E^{\prime}$.
The theme of this Section is that the PFB-POV only loosely links symmetry and geometry: neither one determines the other; both conditions of (5.3) can independently fail. Let us now make this looseness concrete with three perfectly ordinary PFB-POV examples: $U(1)$ acting on ${\mathbb{C}}^{3}$ by scalar multiples of the identity; $SO(4)$ acting on ${\mathbb{C}}^{2}$ via a spinor map; and the trivial action of $SU(n)$ on ${\mathbb{C}}^{m}$. Each case packs a lesson.
@@
From the PFB-POV this is perfectly admissible: one may posit a group that preserves the relevant geometric structures without exhausting them. Still, there is a geometric interpretation: the action rotates the complex volume form of $\mathbb{C}^{3}$. [^29] That is, it rotates $\Lambda^{3}\mathbb{C}^{3}$ (equivalently, the determinant line). Fixing a unit complex volume form reduces $U(3)$ to $SU(3)$; the residual $U(1)$ is the phase on the determinant line. See footnote 23. Thus in this case we have
@@
Lesson: even with a faithful representation, the full geometry of the fibre may not be recovered from $\rho(G)$ alone.
@@
$\rho(G)=\mathbb{1}\subset\mathsf{Aut}(V),\quad\dim(G)\leq\dim(\mathsf{Aut}(V))\ \text{if }m\leq n,\quad\dim(G)\geq\dim(\mathsf{Aut}(V))\ \text{if }m\geq n.$ (5.6)
Even if $n=m$, one cannot reconstruct $G$ from its representation on $V$, since the action is trivial. Thus we have
@@
so both conditions in (5.3) fail. Intrinsically, $G$ may be either larger or smaller than $\mathsf{Aut}(V)$, and the group can neither be recovered from, nor recover, the geometry of the fibre. Lesson: with a trivial representation, the fibre carries no information about $G$, and $G$ imposes no structure on the fibre.
Example 3 (non-faithful but geometrically admissible). Consider $G=SO(4)$ with fibres $V=\mathbb{C}^{2}$. $SO(4)$ admits two inequivalent irreducible representations on $\mathbb{C}^{2}$, corresponding to left- and right-handed spinors under $SU(2)$. If we pick one of these factors to act, $SO(4)$ does preserve the structure of $\mathbb{C}^{2}$, and so the situation is admissible from the PFB-POV. In this case we have
@@
The image of the representation matches $\mathsf{Aut}(V)$, yet the full group $G$ is strictly larger. Lesson: even when $\rho(G)\simeq\mathsf{Aut}(V)$, the embedding of $G$ may lack a clear geometric rationale.
Let us sum up the examples. In the $U(1)$ case on $\mathbb{C}^{3}$ (Eq. (5.5)), the representation was faithful, so $G\simeq\rho(G)$. Here the group can be grounded in appropriate geometric structures on the fibre, but it does not reflect the geometry fully. By contrast, in the cases of $SO(4)$ acting on $\mathbb{C}^{2}$ and $SU(n)$ acting trivially on $\mathbb{C}^{m}$, the representation is not faithful: $G\not\simeq\rho(G)$, so no unique reconstruction of the group from the fibre is possible, even in principle.
What we can abstract from the examples is that the two necessary conditions for the group to reflect the geometry of the fibre, as expressed in Equation (5.3) can fail independently. We may have $G\not\simeq\rho(G)$, or $\rho(G)\subset\mathsf{Aut}(V)$. And even when $G\simeq\mathsf{Aut}(V)$ (add that $\mathsf{Aut}(V)$ acts transitively, if you want), the link between $\mathsf{Aut}(V)$ and GG—namely $\rho(G)$—may still fail. These cases illustrate the PFB-POV’s tolerance for symmetry groups that fall short of reflecting the geometry of the typical fibers of a vector bundle.
@@ 5.2 Fundamental obstacles to equivalence
It is instructive to look at the mechanism behind this in more detail. Since $e^{i(\theta+2\pi)}=e^{i\theta}$, any continuous homomorphism $\rho:U(1)\!\to\!\mathbb{C}^{\times}$ would be constrained by:
$e^{in(\theta+2\pi)}=e^{in\theta}e^{i2\pi n}=e^{in\theta}$ (5.10)
so must satisfy $e^{i2\pi n}=1$. If, however, one replaces $U(1)$ by a non-compact group, such as its universal cover $\mathbb{R}$, the general continuous characters into $\mathbb{C}^{\times}$ are
$\rho_{\lambda}(t)=e^{\lambda t},\qquad\lambda\in\mathbb{C},\quad t\in\mathbb{R}$ (5.11)
so there is a continuum of representations, including varying modulus when $\mathsf{Re}(\lambda)\neq 0$. [^30] But even restricting to $\mathsf{Re}(\lambda)=0$, i.e. to $\lambda=iq$, would still allow $q\in\mathbb{R}$. That is, since we don’t identify $t\equiv t+2\pi$ in the domain of the representation, the image, $e^{iq(t+2\pi)}=e^{iqt}e^{iq2\pi}$ is not necessarily equal to $e^{iqt}$, and so there is no need to identify $e^{iq2\pi}=1$. Thus, in the symmetry-first framework, discreteness arises only from a topological identification built into the choice of group, not from the bare formalism.
In the geometry-first VB-POV the reasoning is inverted. Here one begins not with a Lie group but with a complex line bundle $(E,M,V)$ whose fibre $V\simeq\mathbb{C}$ carries no additional metric or orientation, so that $\mathsf{Aut}(V)\simeq\mathbb{C}^{\times}$. All further fields are are elements of spaces built tensorially from this fundamental bundle, and distinct “charges” correspond to the tensor powers $E^{\otimes n}$ and their duals. The integer label $n$ simply counts how many copies of the fundamental fibre enter the construction; the possible weights therefore form a discrete lattice. Even though, as a group, $\mathsf{Aut}(V)$ admits continuously many characters, the geometry-first formalism can reproduce only those obtained by finite tensor operations. Quantisation of charge thus follows not from topology or compactness but from the internal combinatorics of the tensor algebra. It is also shows how the VB-POV is falsifiable: it cannot accommodate irrational charge ratios.
@@
A different obstacle to equivalence arises with the exceptional Lie groups. The VB-POV proceeds by fixing a fibre $V$ endowed with invariant geometric or algebraic data—an inner product, symplectic or volume form, or some higher-rank tensor—and defining the gauge group as
@@
For the classical families this procedure is canonical: the data determine a unique group whose action both preserves and exhausts the geometry of $V$. For the exceptional families, however, the correspondence between geometry and symmetry becomes tenuous.
[…]
5.3 Composite fibres and the Standard Model
In many gauge theories—and routinely in the Standard Model—the fibres of associated bundles take the form of direct sums or tensor products, such as

### v5

@@ 5.1 Slack between symmetry and geometry: examples and lessons
Recall from Section 2.1 that an associated bundle $E=P\times_{\rho}V$ requires only that $\rho(G)$ preserve the structure of $V$:
@@
The geometry-first picture demands more. It requires
@@
Both conditions can fail independently, and the PFB-POV tolerates all such failures. Three examples make this concrete.
Example 1 (faithful but not surjective). The condition $G\simeq\rho(G)$ holds iff $\rho$ is faithful (injective), so that only $\mathbb{1}$ acts trivially. Consider the gauge group $U(1)$, fibres $\mathbb{C}^{3}$ with Hermitian inner product, and the representation
$\rho(\theta)=e^{i\theta}\mathbb{1}.$
@@
From the PFB-POV this is perfectly admissible: one may posit a group that preserves the relevant geometric structures without exhausting them. Still, there is a geometric interpretation: the action rotates the complex volume form of $\mathbb{C}^{3}$. [^30] That is, it rotates $\Lambda^{3}\mathbb{C}^{3}$ (equivalently, the determinant line). Fixing a unit complex volume form reduces $U(3)$ to $SU(3)$; the residual $U(1)$ is the phase on the determinant line. See footnote 25. Thus in this case we have
@@
$\rho(G)=\mathbb{1}\subset\mathsf{Aut}(V).$ (5.4)
[…]
@@
so both conditions in (5.2) fail. Intrinsically, $G$ may be either larger or smaller than $\mathsf{Aut}(V)$, and the group can neither be recovered from, nor recover, the geometry of the fibre.
Example 3 (non-faithful but geometrically admissible). Consider $G=\mathrm{Spin}(4)\cong SU(2)\times SU(2)$ with fibres $V=\mathbb{C}^{2}$. $\mathrm{Spin}(4)$ admits two inequivalent irreducible representations on $\mathbb{C}^{2}$, corresponding to left- and right-handed chiral spinors. If we pick one of these factors to act, $\mathrm{Spin}(4)$ does preserve the structure of $\mathbb{C}^{2}$, and so the situation is admissible from the PFB-POV. In this case we have
@@
The image of the representation matches $\mathsf{Aut}(V)\simeq SU(2)$, yet the full group $G$ is strictly larger.
The upshot is that the two conditions of (5.2) can fail independently. A faithful representation ($G\simeq\rho(G)$) need not be surjective onto $\mathsf{Aut}(V)$; a surjective image ($\rho(G)\simeq\mathsf{Aut}(V)$) need not faithfully represent the full group. The principal-bundle framework tolerates all of these mismatches. So much for the simple cases.
5.1.a A side-by-side comparison: $SU(n)$ gauge theory
It may help to see both formulations at work on a single example. Take a Hermitian vector bundle $(E\to M,\langle\,\cdot\,,\,\cdot\,\rangle,\epsilon)$ with typical fibre $\mathbb{C}^{n}$, inner product $\langle\,\cdot\,$$\cdot\,\rangle$, and complex volume form $\epsilon$ (i.e. a nowhere-vanishing section of $\Lambda^{n}E$). The automorphism group preserving both structures is $\mathsf{Aut}(E_{x})\simeq SU(n)$.
VB-POV.
One begins with $E$ and a compatible covariant derivative $\nabla$. All further structure is derived:
  • The adjoint bundle is $\mathsf{End}_{0}(E):=\{A\in\mathsf{End}(E)\mid\mathsf{Tr}(A)=0,\,A^{\dagger}=-A\}$, the bundle of traceless anti-Hermitian endomorphisms. Its sections are the “gluon fields” in the sense that $\nabla={\mathrm{d}}+\omega$ for $\omega\in\Gamma(T^{*}M\otimes\mathsf{End}_{0}(E))$.
  • The determinant line bundle is $\det(E):=\Lambda^{n}E$. The volume form $\epsilon$ trivialises it; this is what reduces $U(n)$ to $SU(n)$.
  • Matter multiplets are sections of $E$ (fundamental), $E^{*}$ (anti-fundamental), $\mathrm{Sym}^{k}(E)$ (symmetric tensors), $\Lambda^{k}E$ (antisymmetric), etc. Each inherits a covariant derivative from $\nabla$ by the Leibniz rule.
All of this is determined by $(E,\langle\,\cdot\,,\,\cdot\,\rangle,\epsilon,\nabla)$. No group is postulated; $SU(n)$ appears only when one asks what transformations preserve the data.
PFB-POV.
One begins instead with a principal $G$-bundle $(P,M,G)$ and a principal connection $\varpi$. To recover the same physics, one could take $G=SU(n)$ acting in the fundamental representation $\rho_{\mathbf{n}}$, in which case $E=P\times_{\rho_{\mathbf{n}}}\mathbb{C}^{n}$ and $\nabla$ is induced by $\varpi$ as in (A.6). In this case the two formulations coincide.
But the PFB-POV does not require this choice. One could equally well take:
  1. 1.
    $G=U(n)$ with $\rho=\rho_{\mathbf{n}}$: now $\dim(G)>\dim(\mathsf{Aut}(V))$, and there is an extra $U(1)$ factor that acts trivially on the orientation $\epsilon$—slack in the sense of Example 1.
  2. 2.
    $G=SU(n)\times SU(m)$ with $\rho$ the projection onto the first factor: the second factor acts trivially on $E$, and cannot be recovered from $\mathsf{Aut}(E_{x})$—slack in the sense of Example 2.
  3. 3.
    $G=Spin(2n)$ with $\rho$ a spinor representation onto $\mathbb{C}^{2^{n-1}}$: the map is surjective but not injective, so $G\not\simeq\rho(G)$—slack in the sense of Example 3.
The VB-POV rules out all three by fiat: it insists that the gauge group be the automorphism group of the fundamental bundle. But each of these is a perfectly consistent principal-bundle formulation, and each induces a well-defined covariant derivative on the associated bundle. Yet in each case the group $G$ floats free of the geometry of the fibre: it is either too large, too small, or simply different.
5.1.b The slack world
Consider a world with a single matter field: a section of $E$ with typical fibre $V\simeq(\mathbb{C}^{n},\langle\cdot,\cdot\rangle)$. The automorphism group preserving only the inner product is $\mathsf{Aut}(V)=U(n)$. But suppose the principal bundle has structure group $G=SU(n)$, acting in the fundamental representation. Then $G$ preserves more structure—an orientation—than the matter fields intrinsically possess.
In the PFB-POV, this is perfectly consistent. The connection $\varpi$ carries physical observables independently of matter. The Yang–Mills equations are valued in $\mathfrak{su}(n)$, not $\mathfrak{u}(n)$; the gauge sector has different degrees of freedom. In vacuum Yang–Mills theory, the connection $A_{\mu}$ is a physical field in its own right, with dynamics governed by $G$. If $G=U(n)$ rather than $SU(n)$, the vacuum theory has an extra $U(1)$ factor. These are different theories, with different physical content. Moreover, any Lie-algebra directions in $\ker\rho_{*}$ remain unsourced by matter: the matter current $J(A,\Phi)$ takes values in $\mathrm{im}(\rho_{*})$, so the slack directions satisfy vacuum Yang–Mills equations regardless of the matter configuration (cf. Equation (A.11) in Appendix A).
Call this the slack world: $G$ has physical significance independently of matter, and $G\neq\mathsf{Aut}(V)$. [^31] Parallel transport around a closed curve at $x\in M$ defines an element of $\mathsf{Aut}(E_{x})$. The subgroup generated by all such curves is the holonomy group $\mathsf{Hol}(\nabla)\subseteq\mathsf{Aut}(V)$. Under natural assumptions, a principal bundle $(P^{\prime},M,G^{\prime})$ exists with $G^{\prime}\simeq\mathsf{Hol}(\nabla)$ (cf. \@BBOPcite\@BAP\@BBN(Michor, 2008, Theo. 17.11)\@BBCP). But if the posited $G$ is larger than $\mathsf{Aut}(V)$, the slack persists. Evidence for $G$ comes from the gauge sector; evidence for $\mathsf{Aut}(V)$ from the matter sector. The two need not coincide.
But that is not our world—at least, not our Standard Model; we do not observe such a mismatch. To be clear: this is not a complaint that practitioners of the PFB-POV secretly believe the principal connection $\varpi$ is a physical entity—though most attribute physical significance to its spacetime representatives $A_{\mu}$, or to a section of the affine bundle of connections \@BBOPcitep\@BAP\@BBN(Jacobs, 2023; Kobayaschi, 1957)\@BBCP; even in vacuum Yang–Mills theory where no matter fields are present. The complaint is that the PFB-POV’s representational resources outrun its explanatory needs. This freedom could in principle be observable—as the slack world illustrates—but in the theories physicists actually take seriously, it is not. I return to this point in Section 5.4.
@@ 5.2 Fundamental obstacles to equivalence
In the geometry-first VB-POV the reasoning is inverted. Here one begins not with a Lie group but with a complex line bundle $(E,M,V)$ whose fibre $V\simeq\mathbb{C}$ is equipped with whatever structure one takes to be physically fundamental—a Hermitian inner product if one wishes to preserve probability norms, or nothing beyond the linear structure if not—so that $\mathsf{Aut}(V)$ is $U(1)$ or $\mathbb{C}^{\times}$ accordingly. All further fields are elements of spaces built tensorially from this fundamental bundle, and distinct “charges” correspond to the tensor powers $E^{\otimes n}$ and their duals. The integer label $n$ simply counts how many copies of the fundamental fibre enter the construction; the possible weights therefore form a discrete lattice. Even though, as a group, $\mathsf{Aut}(V)$ may admit continuously many characters, the geometry-first formalism can reproduce only those obtained by finite tensor operations.
The claim is best stated conditionally: if the fundamental charged object is a line bundle $E$ (with whatever fibre structure one stipulates) and all charged fields arise as finite tensor powers and duals of $E$, then the charge ratios lie in a lattice generated by that choice. Irrational ratios cannot arise within such a subtheory—indeed the very phrase ‘irrational tensor power’ is nonsensical.
[…]
I find this explanation considerably more satisfying than the topological one.
@@
A different kind of obstacle arises with the exceptional Lie groups. The issue here is not that the VB-POV’s machinery fails to run—one can always build associated vector bundles given any representation of any group—but that the explanatory direction I have been advertising becomes strained.
The VB-POV proceeds by fixing a fibre $V$ endowed with invariant geometric or algebraic data—an inner product, symplectic or volume form, or some higher-rank tensor—and defining the gauge group as
@@
For the classical families this procedure is canonical: the data determine a unique group whose action both preserves and exhausts the geometry of $V$. For the exceptional families, however, the correspondence between geometry and symmetry becomes less transparent.
Some exceptional groups do admit “matter-first” realisations, but these are formal and not entirely compelling. $G_{2}$ arises as the automorphism group of the octonions, and can be realised as the stabiliser of a generic three-form on $\mathbb{R}^{7}$; $F_{4}$ is the automorphism group of the exceptional Jordan algebra $J_{3}(\mathbb{O})$; $E_{6}$ preserves the cubic norm on the same $27$-dimensional space. But there are two difficulties. First, the structures involved—octonion multiplication, Jordan algebra products, cubic norms—are not “geometric” in the low-level tensorial sense employed elsewhere in this paper (inner products, volume forms, and the operations of the tensor algebra), even though some of them—such as $G_{2}$ as the stabiliser of a three-form—are perfectly geometric in the broader sense of differential geometry. The point is not that these realisations are illegitimate, but that they do not fit the VB-POV’s particular explanatory pattern: start with a fibre endowed with familiar linear-algebraic structure, and read off the group. Second, these realisations are not unique: distinct algebraic structures can yield the same abstract group, so the explanatory arrow from geometry to symmetry is not sharp.
The hard case is $E_{8}$. Its minimal faithful representation is the adjoint, of dimension $248$. One can write down a vector bundle whose automorphism group is $E_{8}$:
$E=(\mathfrak{e}_{8},\kappa,[\cdot,\cdot]),$ (5.9)
where $\mathfrak{e}_{8}$, of dimension $248$, carries the Killing form $\kappa$ and the Lie bracket $\cdot$$\cdot$]—but this is just the Lie algebra $\mathfrak{e}_{8}$ in disguise. There is no small “matter” representation analogous to the fundamental representations of the classical groups. If one tries to run the VB-POV for an $E_{8}$ gauge theory, the “fundamental bundle” is essentially the gauge structure itself rather than a space where matter fields naturally reside. The VB machinery still runs, but the claim of explanatory priority—“start with matter bundles, recover gauge group”—loses its force when the fundamental bundle is already the adjoint.
This is a genuine limitation of the geometry-first rhetoric, though not of the VB formalism per se. It suggests that the VB-POV is most naturally suited to theories whose gauge groups are classical (or at least admit low-dimensional faithful representations), and that $E_{8}$ grand unified theories—if they are to be taken seriously—may require a different philosophical framing. Whether this counts against the VB-POV or against $E_{8}$ unification is a question I leave open.
One might also worry about supersymmetry. A supersymmetric theory requires a symmetry relating bosons to fermions—connecting the gauge-boson sector to matter. Does this not require stepping outside the geometry-first picture? In fact, standard $\mathcal{N}=1$ supersymmetric gauge theories with classical groups fit comfortably within it. The fundamental data are a Hermitian vector bundle $E\to M$ with compatible covariant derivative, together with the spacetime spin bundle $S\to M$. Gauginos are sections of $S\otimes\mathsf{End}_{0}(E)$; matter fermions are sections of $S\otimes E$ or $S\otimes E^{*}$. Supersymmetry transformations mix these fields, but all are sections of bundles constructed tensorially from $(E,\nabla,S)$. Supersymmetry does not require treating the connection as a field of the same type as matter; it requires only that the gaugino—a spinor-valued section of the adjoint bundle—be related to matter fermions. And the adjoint bundle $\mathsf{End}_{0}(E)$ is already part of the tensorial apparatus. Supersymmetry is not a counterexample; it is a stress test that the geometry-first picture passes.
5.3 Recovery of the gauge group from matter bundles
Section 5.1 showed that the gauge group can float free of the geometry of any single matter bundle. But the Standard Model has many particle species, each in its own associated bundle. One might hope that, taken together, they pin down $G$: that the slack disappears when the full matter content is taken into account. If so, positing $G$ independently would be a convenience, not a substantive addition. But the hope fails. To recover the product group from the totality of matter bundles, one must also specify which (factors of which) bundles are fundamental—precisely what the VB-POV supplies at the outset.
In more detail: in many gauge theories—and routinely in the Standard Model—the fibres of associated bundles take the form of direct sums or tensor products, such as